OpenStudy (anonymous):

The inverse of y = 5 - 3x is

OpenStudy (anonymous):

switch the x & y x = 5 - 3y subtract 5 from each side x - 5 = -3y divide by -3 on each side so (x-5)/3 −x+53=y Or 5−x3=y